1 10.00 points 7· At the end of 20X6, Home Ltd. reported the following in shareholders equity: Common shares, no-par value; authorized, unlimited shares issued, 14,590,000 shares Retained earnings s 19,047,000 53,050,000 s 72,097,000 At this time, the shares were trading in the range of $6 to $8 per share on public stock markets. The companys board of directors is contemplating two alternative courses of action: 1. Declaring a 50% stock dividend, or 2. Executing a 3-for-2 stock split Required 1. Prepare the shareholders equity section for each alternative, assuming that market value is used to capitalize the stock dividend. Before (Not Required) Required) Stock Dividend Stock Split Common shares, no-par value; authorized, unlimited shares, issued Retained earnings 2. What would the expected share price be assuming a share price of $4 for alternative 1 and a share price of $6 for alternative 2? (Round your answers to 2 decimal places.) Case 1 Case 2
Answer #1

Total Number of outstanding Shares of Home Ltd. By 2016 with par value $19,047,000 = 14,590,000      
Par value of Share = Total Par value/Total No. of Outstanding Shares = 19,047,000/1590000 = $1.3054      
Retained Earnings = $53,050,000      
      
Let's consider the Average Share price on Public Stock Exchange = $6+$8/2 = $7      
Stock Dividends (on Market value) = 50% of $7 = $3.5      

Before Stock Dividend Stock Split
Common Shares, no par value, Authorized, unlimited shares, issued 14,590,000 = $(14,590,000 x 3.5) = $54,215,000 2/3 x $7 x 14,590,000= $4.67
= $68,086,667
Retained Earnings 53,050,000 53,050,000 $53,050,000

Expected Share Price at share price of $4 at 50% Stock Dividend option = 0.50 x$4 + $4 = $6
Expected Share Price at share price of $6 at 3 to 2 Stock Split option = 2/3 x $6 = $4